If hepatic blood flow is 90 L/h in a 70 kg person and hepatic clearance of morphine is 60 L/h/70 kg, what is the first pass extraction ratio?
Answer #1

Hepatic blood flow = 90 L/h in a 70 kg person

Hepatic clearance of morphine = 60 L /h/70 kg

First pass extraction ratio = Hepatic clearance / Hepatic blood flow

= 60 / 90 = 2/3 = 0.66

It is an intermediate condition since ratio is in the range of (0.3-0.7)